# Linying Lv

### v2

@@ 1 Introduction
Sell-side analysts play a crucial role in financial markets by processing information for investors. Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021). This raises a fundamental question: Do the analysts generate value for their clients, or are they merely peddling expensive noise?
@@
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information beyond simple sentiment dimension and the context has been perceived at least as important as the quantifiable outputs. According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
@@
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024). For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the “earnings analysis” topic, the latter provides additional contextual information and reasoning that can be interpreted and analyzed by LLMs.
@@
I quantify the value of analyst information in analyst reports through three general steps: analyst output representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts from I/B/E/S, such as stock recommendations, earnings forecasts, and target prices. For structured representation of qualitative textual information, I extract contextualized representations of analyst report texts using state-of-the-art LLMs. Each report is mapped into a structured embedding space using the LLMs’ transformer architecture.
@@
With structured text representation, I evaluate how well analyst information explains stock returns using an R-squared measure, which also serves as the terminology of information content. A significant challenge in projecting stock returns onto analyst outputs is the risk of in-sample overfitting due to the high-dimensionality of text embeddings. To address this issue, I employ Ridge regression, which introduces a regularization term to penalize large coefficients. This regularization helps in handling the high-dimensional embeddings while preserving the model’s ability to capture essential relationships between analyst outputs and stock returns. Additionally, I evaluate the out-of-sample explanatory power of quantitative and qualitative measures for abnormal stock returns to assess model stability.
@@
I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using each type of information individually. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content within analyst report text.
The informativeness of analyst reports varies considerably depending on whether they are associated with forecast revisions or reiterations. Both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions. Specifically, when analyst reports involve recommendation revisions, the combined use of numerical and textual data yields an out-of-sample $R^{2}$ of 22.63%. In stark contrast, for reports with earnings forecast reiterations, the out-of-sample $R^{2}$ is not statistically significant, suggesting that such reports contribute minimal incremental information to the market.
@@
A notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al. (2014), Kim and Song (2015), Lobo et al. (2017), and Barron et al. (2017), An investigation on samples within and beyond the earnings announcement windows reveals that the information content of analyst reports peaks in the first week following earnings announcements. The finding supports Huang et al. (2018)’s perspective that timely reactions to corporate disclosures offer valuable information advantages to clients. Using separately trained models on samples within and beyond earnings announcement windows, the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods. The significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
Analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models to isolate the unique information from analyst reports. The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights.
@@
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains an open question. To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results show that over half of the explanatory power of the report, in terms of out-of-sample $R^{2}$, is attributed to Income Statement Analyses in analyst reports, particularly the interpretation of realized income numbers. This evidence again emphasizes the critical role analysts play in interpreting financial data and providing the stories and context needed for making investment decisions.
@@
So far, I have measured information value statistically with R-squared. To translate it into economic terms, I adopt Kadan and Manela (forthcoming)’s framework to quantify the dollar value of analyst information to strategic investors. The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts. This multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to the price impact, representing the cost associated with trading on this information.
@@
The findings reveal significant value in analyst information. For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports are \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million from both combined. The annualized aggregate information value is estimated at \$6.89 million, based on an average of 15 days with reports released per year. According to Kadan and Manela (forthcoming), this estimate represents a conservative lower bound. Notably, these values increase for large stocks, bold analysts, and reports released promptly following earnings announcements.
@@
My paper relates to three strands of literature. This paper first relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Prior research often relies on manually encoding textual information and sentiment analyses, which is labor-intensive and may overlook nuanced insights. My work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations. A direct comparison with the Naive Bayes tone classification used in Huang et al. (2014) reveals two main advantages of LLMs: they more effectively capture overall sentiment in analyst reports and extract richer information beyond tone. Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure from 3% to above 10%. The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information. Topic decomposition analysis offers a new dimension for understanding analyst information value by revealing their role in interpreting financial data.
@@
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2024), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in financial contexts. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. This research leverages the additive nature of LLM embeddings and proposes a Shapley value decomposition approach for interpretability, offering a new framework for explainable AI.
@@
Third, this paper deepens the understanding of the economics of tipping - a well-documented phenomenon between analysts and institutional investors (Irvine et al., 2007; Christophe et al., 2010). Tipping involves brokerage firms rewarding high-commission clients with early research access, which allows these firms to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends prior work by quantifying the economic dollar value of such insider information.
@@
The rest of the paper is organized as follows. Section 2 introduces methodology, measures of analyst information value, and interpretation. Section 3 summarizes data construction and empirical results. Section 4 concludes.
@@ 2 Methodology
In this section, I introduce how I employ large language models for topic modeling and text embeddings, the information content estimation, the strategic information value estimation, and the decomposition metric used to evaluate the information value of topics. [^1] The details of machine learning models used to explain contemporaneous stock returns are discussed in the Online Appendix C.
@@ 2.1 Large Language Models
I harness the power of advanced generative AI models to conduct comprehensive analyses of analyst reports. The methodology employs a multi-faceted approach: fine-tuned BERT models for precise sentence classification based on predefined criteria, ChatGPT prompting for extracting and synthesizing major topics discussed by analysts, and LLaMA embedding for representing entire reports as structured high-dimensional vectors.
@@ 2.1.1 ChatGPT Prompting
It ends up with 13 categories defined as follows:
  • Executive Summary: Provides a high-level overview of the report’s key findings and conclusions; includes a brief description of the company, its industry, and the purpose of the report; highlights the most important points from the analysis, such as the company’s financial performance, competitive position, and growth prospects.
  • Company Overview: Offers a comprehensive description of the company, including its history, business model, and key products or services; discusses the company’s organizational structure, management team, and corporate governance; analyzes the company’s mission, vision, and strategic objectives.
  • Industry Analysis: Provides an in-depth analysis of the industry in which the company operates; includes information on market size, growth trends, and key drivers; discusses the regulatory environment, technological advancements, and other external factors affecting the industry; analyzes the industry’s competitive dynamics and the company’s position within the industry.
  • Competitive Landscape: Identifies the company’s main competitors and their market share; compares the company’s products, services, and pricing strategies with those of its competitors; analyzes the strengths and weaknesses of the company and its competitors; discusses potential new entrants and substitutes that could disrupt the competitive landscape.
  • Financial Analysis: Analyzes the company’s revenue, expenses, profitability, assets, liabilities, shareholders’ equity, liquidity, solvency, capital structure, cash flows, and key financial ratios; compares the company’s financial performance with that of its competitors and industry benchmarks; discusses trends in the company’s financial performance over time.
  • Business Segments: Provides a detailed analysis of the company’s various business segments or divisions; discusses the financial performance, growth prospects, and challenges of each segment; analyzes the contribution of each segment to the company’s overall revenue and profitability.
  • Growth Strategies: Discusses the company’s strategies for driving future growth, such as organic growth initiatives, product innovations, and geographic expansions; analyzes the company’s mergers and acquisitions (M&A) strategy and potential targets; examines the company’s investments in research and development (R&D) and marketing.
  • Risk Factors: Identifies and analyzes the key risks facing the company, such as market risks, operational risks, financial risks, and legal/regulatory risks; discusses the potential impact of these risks on the company’s financial performance and growth prospects; examines the company’s risk management strategies and mitigation measures.
  • Management and Governance: Provides an overview of the company’s management team, including their experience, expertise, and track record; analyzes the company’s corporate governance practices, such as board composition, executive compensation, and shareholder rights; discusses the company’s succession planning and key person risks.
  • Environmental, Social, and Governance (ESG) Factors: Analyzes the company’s performance and initiatives related to environmental sustainability, social responsibility, and corporate governance; discusses the potential impact of ESG factors on the company’s reputation, risk profile, and financial performance; examines the company’s compliance with ESG regulations and industry standards.
  • Valuation: Estimates the intrinsic value of the company’s shares using various valuation methodologies, such as discounted cash flow (DCF) analysis, relative valuation multiples, and sum-of-the-parts analysis; compares the company’s valuation with that of its peers and historical benchmarks; discusses the key assumptions and sensitivities underlying the valuation analyses.
  • Investment Thesis: Summarizes the key reasons for investing (or not investing) in the company’s shares; discusses the potential catalysts and risks that could impact the company’s valuation and stock price performance; provides a target price or price range for the company’s shares based on the valuation analyses and investment thesis.
  • Appendices and Disclosures: Includes additional supporting materials, such as financial statements, ratio calculations, and detailed segment data; provides important disclosures, such as the analyst’s rating system, potential conflicts of interest, and disclaimers; discusses the limitations and uncertainties of the analysis and the need for further due diligence by investors.
The Financial Analyses category is further divided into four subcategories: Income Statement Analyses, Balance Sheet Analyses, Cash Flow Analyses, and Financial Ratios. Additionally, a category labeled “None of the Above” is included to accommodate sentences that do not fit into any of the predefined topics. In total, this process generates 17 distinct topics for classification. The distribution of topics is relatively stable across years, as illustrated in Figure 1. To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category. Furthermore, Table LABEL:tab:_topic_example offers illustrative sentences extracted directly from analyst reports, providing concrete examples of the discussions typical to each category.
@@
This training sample serves as the training sample for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. The scarcity of Executive Summary content may result from the difficulty of identifying this category at the sentence level. Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in ESG discussions.
@@ 2.1.2 Text Representations
A systematic textual analysis requires a structured representation of analyst reports. Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. Within LLMs, tokens are mapped to unique identifiers (IDs) based on a predefined vocabulary. This vocabulary is learned during the model’s training phase and contains all the tokens the model can understand.
@@
At the heart of LLMs lies the transformer architecture. The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance. A core output of the transformer architecture is text embeddings. These embeddings are dense, continuous vectors that capture semantic and syntactic information. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
To obtain structured representations of analyst reports, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens. The tokenizer processes the text data, converting it into tokens that the model can understand. By processing the analyst reports through this model, I obtained rich, contextual embeddings that capture the nuanced information within the reports. These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and economic modeling.

### v3

@@ 1 Introduction
Sell-side analysts play a crucial role in financial markets by producing and processing information for investors. Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021). This raises a fundamental question: Do the analysts generate value for their clients, or are they merely peddling expensive noise?
@@
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information beyond the simple sentiment dimension. The context has been perceived at least as important as the quantifiable outputs. According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
@@
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024). Consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences analyze earnings in a negative tone, the latter provides additional contextual information and reasoning. The question now becomes: do investors disregard obvious facts and figures in favor of details in lines?
@@
I investigate the question through three general steps: analyst output representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts like stock recommendations, earnings forecasts, and target prices. For qualitative textual information, I extract contextualized representations by mapping each report into a structured LLM embedding space.
@@
With structured representation, I evaluate how well analyst information explains stock returns as evidence of information content. A significant challenge in projecting stock returns onto text embeddings is the risk of in-sample overfitting due to high dimensionality. To address this issue, I employ Ridge regressions. The simple model penalizes large coefficients while still preserving the model’s ability to capture essential signals that move the market. Additionally, I use out-of-sample R-squared to evaluate model performances.
@@
I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR_{[-1,+1]}$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using either information type alone. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the distinct and economically meaningful information content embedded in analyst report narratives.
The informativeness of analyst reports varies substantially with forecast changes. Analyst reports accompanied by forecast revisions provide significantly more valuable information to financial markets compared to reiterations. For reports containing recommendation revisions, the combination of numerical and textual information generates an out-of-sample $R^{2}$ of 22.63%. In contrast, reports merely reiterating previous earnings forecasts yield statistically insignificant out-of-sample $R^{2}$ is not statistically significant, indicating they provide minimal new information to market participants.
@@
Previous research emphasizes analysts’ role in information processing around corporate earnings announcements( e.g., Livnat and Zhang, 2012, Keskek et al., 2014, Kim and Song, 2015, Lobo et al., 2017, and Barron et al., 2017). My analysis on samples during and outside earnings announcement periods reveals that analyst reports’ information content peaks in the first week following earnings releases. The finding aligns with Huang et al. (2018), who argue that prompt responses to corporate disclosures provide valuable insights to clients. Using distinct models trained on samples during and outside earnings announcement windows, the out-of-sample $R^{2}$ approximately doubles in the first week following earnings announcement dates compared to other periods. The substantial increase highlights the incremental value analysts add by interpreting and contextualizing earnings information for investors.
To address the concern that my results merely capture analysts’ “piggybacking” on earnings conference call information or post-earning-announcement drift (PEAD), I control for both earnings surprises and the latest earnings conference call transcripts using text embeddings. The results show that analyst reports provide distinct and valuable insights even after accounting for these alternative explanations.
@@
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains largely unsolved. I develop a systematic approach to quantify the importance of different report content. By exploiting the additive property of text embeddings and implementing a Shapley value decomposition, I attribute the explanatory power to 17 major topics discussed in analyst reports.
To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results reveal that Income Statement Analyses, particularly interpretations of realized earnings, account for over half of the reports’ explanatory power in terms of out-of-sample $R^{2}$. These results reinforce the critical role of analysts in interpreting financial data and constructing the narratives and context essential for investment decisions.
@@
To translate statistical measures of information content into economic terms, I adopt Kadan and Manela (2024)’s framework to quantify the dollar value of analyst information for strategic investors. The multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination. The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to price impact, representing the cost of trading on this information.
@@
The analysis reveals substantial economic value in analyst information. For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports amount to \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million when combined. Based on an average of 15 report days annually, the aggregate annualized information value reaches \$6.89 million—a conservative lower bound according to Kadan and Manela (2024). This value increases for large-cap stocks, bold analysts, and reports released promptly following earnings announcements.
@@
My paper relates to three strands of literature. It first connects to recent studies of market reactions to security analyst reports (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Earlier investigations primarily rely on manual text coding and sentiment analysis, which are labor-intensive and may overlook nuanced insights. My work demonstrates the superiority of state-of-the-art LLMs in generating contextualized textual representations. Compared to Huang et al. (2014)’s Naive Bayes tone classification, LLMs offer two main advantages: better capture of overall report sentiment and extraction of richer information beyond tone. Moving from simple tone analysis to comprehensive LLM embeddings boosts the $R^{2}$ from 3% to over 10%. The substantial improvement in explanatory power highlights LLMs’ potential to transform textual analysis. Additionally, the topic decomposition approach provides novel insights into analysts’ value by revealing their role in interpreting financial data.
@@
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2024), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in a financial paradigm. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. By exploiting the additive property of LLM embeddings and developing a Shapley value decomposition approach, my research provides a new framework for explainable AI.
@@
Third, this paper advances the understanding of analyst tipping - the practice of brokerage firms providing early research access to high-commission clients (Irvine et al., 2007; Christophe et al., 2010). Tipping allows brokerages to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends by quantifying the economic dollar value of such privileged information.
@@
The rest of the paper is organized as follows. Section 2 introduces methodology, measures of analyst information value, and interpretation framework. Section 3 describes the data and empirical results. Section 4 concludes.
@@ 2 Methodology
This section describes the key methodologies: the use of large language models for topic modeling and text embeddings, estimation of information content and strategic information value, and decomposition metrics for evaluating topic-specific information value. [^1] The details of machine learning models used to explain contemporaneous stock returns are provided in the Online Appendix C.
@@ 2.1 Large Language Models
The analysis of analyst reports leverages three advanced generative AI models. A fine-tuned BERT model classifies sentences according to predefined criteria, ChatGPT extracts and synthesizes major topics through prompting, and LLaMA generates high-dimensional vector embeddings to represent report content.
@@ 2.1.1 ChatGPT Prompting
It ends up with 16 categories defined in Table LABEL:tab:_topic. A category labeled “None of the Above” is included to accommodate sentences that do not fit into predefined topics. To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category. Furthermore, Table LABEL:tab:_topic_example offers illustrative sentences extracted directly from analyst reports, providing concrete examples of the discussions typical to each category.
@@
This labeled sample is used to fine-tune a BERT model, which then classifies all sentences in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. The distribution of topics is relatively stable across years, as illustrated in Figure 1. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least frequent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. Executive Summary sections appear infrequently because they are difficult to identify at the sentence level. While ESG discussions are minimal overall, they increase notably after 2020, reflecting growing market interest in sustainability issues.
@@ 2.1.2 Text Representations
I need a structured representation of writtrn reports. Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. These tokens are then mapped to unique identifiers based on a predefined vocabulary learned during model training.
@@
At the heart of LLMs lies the transformer architecture. A core output of the transformer is text embeddings - dense, continuous vectors that capture semantic and syntactic information. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
For this analysis, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens. The resulting contextual embeddings capture nuanced report information and are suitable for downstream tasks like sentiment analysis and economic modeling.